# Physics Encoded Blocks in Residual Neural Network Architectures for Digital Twin Models

### v1

@@ 5.1.2 Results
The results illustrate the effectiveness of the residual blocks in our general Physics Encoded Neural Network framework in its capability to facilitate the optimization of loss functions when physics models are directly baked into neural architectures. In the next section, we evaluate our general framework on a complex, real-world problem.
[…]
Figure 11: Comparison of DeLaN and PERNN (our method) in terms of closeness to ground-truth. Both models can not only learn the desired torques but also disambiguate the individual torque components even though they were trained on the super-imposed torques. It can be seen that the PERNN model generalizes significantly better than the original DeLaN model and presents better accuracy on all measured variables.
Table 3: Numerical results of the experiments comparing the performance of PERNN against DeLaN on 2-dof robotic motion
Model Torque (MSE) Inertia (MSE) Coriolis & Centrifugal (MSE) Gravitational (MSE)
DeLaN 0.2073 0.41 0.097 0.0223
PERNN 0.00495 0.00458 0.0011 0.0082
@@ 6 Conclusion and Future Work
This paper presents a novel physics-encoded neural network architecture for embedding physics models into neural networks. The concept of residual knowledge blocks allows this architecture to be used as a general framework for encoding physics knowledge into neural networks using any differentiable physics model. It is a promising research direction in terms of adding reliability and interpretability to neural network-based learning algorithms by providing the capability to discover unknown variables in the environment which are not part of the originally observed data. Furthermore, this method provides a direct and effective way to build prototype models rapidly for the general digital twin scenarios with limited data and physical definitions of real-world entities.
We investigated our method in two distinct applications. The first involved modelling basic robotic motion using Euler-Lagrangian equations of motion as physics prior. We extended the current state-of-the-art method (DeLaN) to build our residual knowledge blocks and presented the effectiveness of the residual methodology in the context of our general physics-encoded neural network framework. The second application involved developing a steering model for autonomous vehicles in the TORCS simulation environment. We presented some key achievements here, including significantly improved generalizability of the driving model to new road scenarios compared to pure data-driven, fully-connected neural networks and the well-known physics informed (regularized) neural network method. This was achieved with much less model complexity and data requirements. We showed that this method outperformed both counterparts on unseen test track scenarios regarding driving metrics like average distance travelled before boundary hit and smoothness of steering trajectory. We presented an analysis of the effectiveness of our method to discover unknown variables such as the lookahead distance and the heading difference to the selected point on the road and how the model intelligently predicted these variables based on different road scenarios. This also presented valuable insight regarding the explainability of the model’s decision-making based on what point on the road is selected to be followed for different scenarios.
It would be promising to test this approach further on a wide range of problems - specifically in the digital twin space where data availability is constrained and physics models are ‘ab-inito’ or first principle. Further study on the behaviour of gradient flow through the network could help develop more robust residual blocks, especially for more complex physics where large physics blocks could present more constraints on the algorithm’s convergence.
@@
  • Authors’ contributions: Muhammad Saad Zia and Ashiq Anjum have equal contributions to this manuscript. Lu Liu is an academic reviewer and collaborator on the work in this manuscript. Anthony Conway and Anasol Peña-Rios are industrial collaborators on the work in this manuscript.

### v2

@@ 4.2.1 Flux Tower Data
H $\text{W m}^{-2}$ Sensible heat flux density.
Tau $\text{kg m}^{-1}\text{s}^{-2}$ Momentum flux.
RH % Relative humidity at 2 m.
VPD HPa Vapor pressure deficit.
$\text{R}_{\text{g}}$ $\text{W m}^{-2}$ Global radiation.
Ustar $\text{m s}^{-1}$ Friction velocity.
Tsoil1 ${}^{\circ}\text{C}$ Soil temperature at a depth of 0.05 m.
Tsoil2 ${}^{\circ}\text{C}$ Soil temperature at a depth of 0.05 m.
$\text{T}_{\text{air}}$ ${}^{\circ}\text{C}$ Air temperature at 2 m.
4.2.2 The Net Ecosystem Exchange Model
The Net Ecosystem Exchange (NEE) quantifies the net flux of CO2 between an ecosystem and the atmosphere. It is defined as the balance between ecosystem respiration and photosynthetic uptake. Specifically, NEE can be expressed as:
$\text{NEE}_{t}=\text{R}_{\text{eco},t}-\text{GPP}_{t},$ (3)
where $\text{R}_{\text{eco},t}$ represents the rate of CO2 release due to biological respiration, and $\text{GPP}_{t}$ denotes the Gross Primary Production associated with photosynthesis. By convention, negative NEE values indicate net CO2 uptake by the ecosystem [35, 41].
For periods when global radiation is insufficient (i.e., $\text{R}_{\text{g}}<20\,\text{W.m}^{-2}$), photosynthesis is assumed to be negligible, resulting in $\text{GPP}_{t}\approx 0$. Under these conditions, the measured NEE predominantly reflects ecosystem respiration. The temperature dependence of ecosystem respiration is commonly modelled using an Arrhenius-type expression. Accordingly, the respiration term at night is formulated as:
$\text{R}_{\text{eco},t}=r_{night}\exp\!\Biggl{(}E_{0}\Bigl{(}\frac{1}{T_{ref}-T_{0}}-\frac{1}{\text{T}_{\text{air},t}-T_{0}}\Bigr{)}\Biggr{)},$ (4)
where $r_{night}$ (in $\upmu\text{mol C m}^{-2}\text{s}^{-1}$) is the baseline respiration rate at the reference temperature $T_{ref}=15^{\circ}\text{C}$, $E_{0}$ (in ${}^{\circ}\text{C}$) is a constant that characterizes temperature sensitivity, $\text{T}_{\text{air},t}$ is the ambient air temperature at time $t$, and $T_{0}$ is a fixed offset (typically $-46.02^{\circ}\text{C}$) [35]. For consistency, the parameter $r_{night}$ is periodically updated—commonly every five days—using estimations derived from 15-day windows of historical observations.
This formulation of NEE as described above forms the foundation of our modelling approach.
4.2.3 ODE-based Net Ecosystem Exchange Model
In our approach, the dynamics of Net Ecosystem Exchange (NEE) are modeled using an ordinary differential equation (ODE) that captures the temperature-dependent behaviour of ecosystem respiration. Specifically, the evolution of NEE over time is expressed as
$\frac{d\,\text{NEE}_{t}}{dt}=\frac{d}{d\,\text{T}_{\text{air},t}}\text{R}_{\text{eco},t}(\text{T}_{\text{air},t})\,\frac{d\,\text{T}_{\text{air},t}}{dt},$ (5)
where $\text{R}_{\text{eco},t}(\text{T}_{\text{air},t})$ denotes the ecosystem respiration rate, modeled via an Arrhenius-type function given in Equation 4.
The derivative of the respiration function with respect to air temperature is given by
$\frac{d}{d\,\text{T}_{\text{air},t}}\text{R}_{\text{eco},t}(\text{T}_{\text{air},t})=\frac{E_{0}}{(\text{T}_{\text{air},t}-T_{0})^{2}}\,\text{R}_{\text{eco},t},$ (6)
and the rate of change of the ambient air temperature is modeled as
$\frac{d\,\text{T}_{\text{air},t}}{dt}=\pi\,\frac{\Delta\text{T}_{\text{air},t}}{t_{day}}\,\sin\!\Bigl{(}2\pi\,\frac{t-t_{T_{\max}}}{t_{day}}\Bigr{)}.$ (7)
By incorporating Equation 5 into our Physics Encoded Residual Neural Network (PERNN) framework, we directly embed the known temperature-driven dynamics of ecosystem respiration into the digital twin. This integration ensures that while the neural network components can infer latent variables and correct for discrepancies, the overall model remains grounded in the established physical relationships governing CO2 exchange.
4.2.4 Physics Encoded Residual Neural Network for NEE ODE
In our NEE modeling application, we extend the Physics Encoded Residual Neural Network (PERNN) framework to forecast the evolution of Net Ecosystem Exchange (NEE) by integrating the known deterministic dynamics of ecosystem respiration into a digital twin. The goal is to predict the rate of change of NEE, $\frac{d\,\text{NEE}}{dt}$, and subsequently obtain the forecasted NEE at the next time step via
$\text{NEE}_{t+1}=\text{NEE}_{t}+\widehat{\frac{d\,\text{NEE}}{dt}}\cdot\Delta t,$ (8)
where $\Delta t$ is 30 minutes, consistent with the measurement interval of the flux data.
Let $\mathbf{x}\in\mathbb{R}^{d}$ denote the vector of input features, where $d$ corresponds to the number of variables defined in Table 1 (e.g., $\text{T}_{\text{air}}$, $\text{R}_{\text{g}}$, RH, VPD, etc.). These features encapsulate the meteorological and temporal context required for the NEE model.
The learning block, denoted by $\mathcal{L}(\mathbf{x},\phi_{learn})$, is implemented as a set of fully connected layers. It processes the input state $\mathbf{x}$ and infers the intermediate variables that are not directly observable:
$\mathcal{L}(\mathbf{x},\phi_{learn})=\begin{bmatrix}E_{0}\\ r_{b\_night}\\ \left(\frac{dT}{dt}\right)_{pred}\end{bmatrix}.$
Here, $E_{0}$ represents the temperature sensitivity, $r_{b\_night}$ is the base respiration at the reference temperature, and $\left(\frac{dT}{dt}\right)_{pred}$ is the predicted rate of change of ambient air temperature. These variables are subsequently fed to the physics block.
The physics block, $\mathcal{P}(\mathbf{x},\mathcal{L})$, is a non-trainable computational graph that encapsulates the physical operators derived from the ODE governing NEE dynamics. Using the predicted temperature derivative $\left(\frac{dT}{dt}\right)_{pred}$ from the learning block, the physics block computes the change in NEE based on Equation 10 as:
$\mathcal{P}(\mathbf{x},\mathcal{L})=\frac{d\,\text{NEE}_{t}}{dt}=\frac{E_{0}}{(\text{T}_{\text{air},t}-T_{0})^{2}}\,\text{R}_{\text{eco},t}\,\left(\frac{dT}{dt}\right)_{pred}.$ (9)
To compensate for any discrepancies between the physics-based prediction and the actual dynamics observed in the data, a residual block $\mathcal{R}(V;\phi_{res})$ is employed. Here, $V$ represents an intermediate feature vector extracted from the learning block. The residual block predicts an additive correction $r$ such that the final prediction of the rate of change in NEE is:
$\widehat{\frac{d\,\text{NEE}}{dt}}=\mathcal{P}(\mathbf{x},\mathcal{L})+r.$ (10)
Ground-Truth Computation
Ground-truth values for the rates of change are computed from the flux tower measurements using a right-sided first-order finite difference approximation. Specifically, for NEE:
$\frac{d\,\text{NEE}_{t}}{dt}\approx\frac{\text{NEE}_{t+1}-\text{NEE}_{t}}{\Delta t},$ (11)
and similarly for air temperature:
$\frac{d\,T_{air,t}}{dt}\approx\frac{T_{air,t+1}-T_{air,t}}{\Delta t},$ (12)
with $\Delta t=30$ minutes.
Parameter Estimation for $E_{0}$ and $r_{night}$
Since the flux tower data does not directly provide ground-truth values for the parameters $E_{0}$ and $r_{night}$, these are estimated using the REddyProc partitioning algorithm [42]. REddyProc applies a non-linear regression approach based on the methodology of [33] to partition the flux data, particularly during nighttime conditions when photosynthetic uptake is negligible. In this context, the Lloyd-and-Taylor model [36] is fitted to the scatter of NEE and $\text{T}_{\text{air},t}$ values, providing estimates within established ranges (e.g., $E_{0}\in[50,400]$ and $r_{night}>0$) as discussed in [35].
Loss Function and Two-Phased Training Procedure
The model is trained to minimize the Mean Squared Error (MSE) loss, with two distinct phases. We use the Mini-Batch Gradient Descent algorithm and Adam Optimizer to optimize the loss function, where $n$ is any arbitrary batch size for each sample in the gradient descent algorithm.
  • Phase I: Warm Start: In this phase, the learning block $\mathcal{L}$ is pre-trained using heuristic labels generated by inverting Equations 11 and 12. The MSE loss is computed solely on the intermediate variables:
    $L_{1}=\frac{1}{n}\sum_{i=1}^{n}\Biggl{[}\Bigl{(}\widehat{E_{0}}^{(i)}-E_{0}^{(i)}\Bigr{)}^{2}+\Bigl{(}\widehat{r}_{b\_night}^{(i)}-r_{b\_night}^{(i)}\Bigr{)}^{2}+\Bigl{(}\widehat{\left(\frac{dT}{dt}\right)}^{(i)}-\left(\frac{dT}{dt}\right)^{(i)}\Bigr{)}^{2}\Biggr{]},$ (13)
    where $n$ is the number of samples in the mini-batch.
  • Phase II: Integrated Training: In the second phase, the full model $M$ — comprising the learning block $\mathcal{L}$, the fixed physics block $\mathcal{P}$, and the residual block $\mathcal{R}$ — is trained end-to-end. The MSE loss in this phase is computed on both the intermediate variables and the forecasted $\text{NEE}_{t+1}$:
    $\begin{split}L_{2}&=\frac{1}{N}\sum_{i=1}^{N}\Biggl{\{}\Bigl{(}\widehat{\text{NEE}}_{t+1}^{(i)}-\text{NEE}_{t+1}^{(i)}\Bigr{)}^{2}\\ &\quad+\Bigl{(}\widehat{E_{0}}^{(i)}-E_{0}^{(i)}\Bigr{)}^{2}\\ &\quad+\Bigl{(}\widehat{r}_{b\_night}^{(i)}-r_{b\_night}^{(i)}\Bigr{)}^{2}\\ &\quad+\Bigl{(}\widehat{\left(\frac{dT}{dt}\right)}^{(i)}-\left(\frac{dT}{dt}\right)^{(i)}\Bigr{)}^{2}\Biggr{\}}.\end{split}$ (14)
    where $\widehat{\text{NEE}}_{t+1}^{(i)}$ is the forecasted value computed via Equation 8.
Figure 9 illustrates the overall PERNN architecture for modeling the NEE ODE, and Algorithm 2 details the two-phased training procedure.
Figure 9: Physics Encoded Neural Network architecture for NEE forecasting. The learning block predicts three intermediate variables: $E_{0}$ (temperature sensitivity), $rb_{night}$ (base respiration) and $dT_{air}$ (change in air temperature). These variables are passed as input to the physics block based on the NEE ODE to predict $d\text{NEE}$ (change in NEE). The intermediate layer output from the learning block is passed to the residual block to predict the residual value to be added to $d\text{NEE}$ for the final prediction of the time derivative $\widehat{d\text{NEE}}$. This is then added to the current $\text{NEE}_{t}$ to output $\widehat{\text{NEE}}_{t+1}$.
Algorithm 2 Two-Phased Training for PERNN for NEE
1:Input: Training dataset $\mathcal{D}$, Loss threshold $\ell_{\text{thresh}}$
2:Output: Trained PERNN model $M$
3:Initialize model $M$ with:
  • Learning block $\mathcal{L}$ (learnable)
  • Physics block $\mathcal{P}$ (non-learnable)
  • Residual block $\mathcal{R}$ (learnable)
4:
5:Phase I: Pre-train Learning Block
6:repeat
7:     Sample a mini-batch from $\mathcal{D}$
8:     Forward pass through $\mathcal{L}$ to predict intermediate variables: $\widehat{E_{0}}$, $\widehat{r}_{b\_night}$, $\widehat{\left(\frac{dT}{dt}\right)}$
9:     Compute MSE loss $L_{1}$ on these variables:
$L_{1}=\text{MSE}\Bigl{(}\begin{bmatrix}\widehat{E_{0}}\\ \widehat{r}_{b\_night}\\ \widehat{\left(\frac{dT}{dt}\right)}\end{bmatrix},\begin{bmatrix}E_{0}\\ r_{b\_night}\\ \left(\frac{dT}{dt}\right)\end{bmatrix}\Bigr{)}$
10:     Backpropagate and update weights in $\mathcal{L}$
11:until $L_{1}$ converges
12:
13:Phase II: Train Integrated Model
14:repeat
15:     Sample a mini-batch from $\mathcal{D}$
16:     Forward pass through $\mathcal{L}$, then through $\mathcal{P}$ (fixed), and finally through $\mathcal{R}$ to obtain forecasted $\widehat{\text{NEE}}_{t+1}$
17:     Compute MSE loss $L_{2}$ on both the intermediate variables and $\widehat{\text{NEE}}_{t+1}$:
$L_{2}=\text{MSE}\Biggl{(}\begin{bmatrix}\widehat{E_{0}}\\ \widehat{r}_{b\_night}\\ \widehat{\left(\frac{dT}{dt}\right)}\\ \widehat{\text{NEE}}_{t+1}\end{bmatrix},\begin{bmatrix}E_{0}\\ r_{b\_night}\\ \left(\frac{dT}{dt}\right)\\ \text{NEE}_{t+1}\end{bmatrix}\Biggr{)}$
18:     Backpropagate and update weights in $\mathcal{L}$ and $\mathcal{R}$
19:until $L_{2}<\ell_{\text{thresh}}$
20:
21:return Trained model $M$
5 Experiments
In this section, we present a comprehensive set of experiments and analyses that build upon the theoretical foundations described in the preceding sections. For the simulation-based autonomous vehicle steering dynamics use case—where the available prior physics is simple and only partially known—we evaluate the performance of our approach in terms of generalizability, data efficiency, and model complexity. In the climate digital twin application, we assess our method’s ability to accurately model Net Ecosystem Exchange (NEE) behavior over time and effectively perform gap-filling, comparing its performance against state-of-the-art techniques. These experiments serve to validate the advantages of our approach in both controlled simulation environments and real-world scenarios.
@@ 5.1.4 Analysis of Unobserved Intermediate Variables
5.2 Net Ecosystem Exchange Modeling
In this section, we present details on the experimentation and results on the problem of NEE modeling for gap-filling in flux-tower observational data. We first present a short analysis on the difference in NEE behavior reflected in the simplistic physics model of NEE based on Equation 4 and the flux-tower measurements. In addition, similar to section 5.1, we present and compare results for PERNN with PENN (Physics Encoded Neural Network) as a variant of our method without the residual block and PINN (Physics Informed Neural Network) that uses the NEE ODE as a regularization on the loss term only, and finally with a conventioanl FCNN (Fully Connected Neural Network). We also present comparison against Random Forest which serves as the current state-of-the-art method for NEE gap-filling [32] and XgBoost, another decision-tree based algorithm similar to Random Forest for extended comparison.
5.2.1 Analysis of NEE Physics Model
Since the NEE ODE (based Equation 10) incorporated in our Physics Block is based on the equation modeled using an Arrhenius-type expression from Equation 4, we compare the NEE dynamics based on this equation with the ground-truth NEE values from the flux-tower observational data. Figure 16 presents four plots from randomly sampled daily cases of NEE observational data. The blue lines represent these observed values, while the red lines represent the values from the NEE physics model calculated using corresponding observed parameter, $T_{air}$ and estimated parameters $E_{0}$ and $r_{b,night}$. It can be observed that the underlying NEE physics model (and consequently the NEE ODE) is an over-simplification of the actual NEE trends in the observational data, albeit following the general progression most of the time.
Figure 16: The comparison of daily NEE values from flux-data and the output from NEE physics model based on Equation 4 for four randomly sampled daily cases from the test-set. The blue line represents the ground-truth observations and the red line represents the values from the NEE physics model. The top row shows samples from dates 2018-08-18 and 2019-10-24, while the bottom row shows samples from dates 2019-03-04 and 2019-08-21 (left to right).
5.2.2 Experimental Setup
We split the flux dataset into training and test sets based on the year of recording. The training set comprised years 2012-2017 (inclusive) while the test set was reserved for years 2018 and 2019 for experimental consistency when measuring gap-filling accuracy over different time scales: daily, weekly, monthly, and quarterly.
For all neural network based methods in this study, we keep the core network architecture and hyperparameters same for experimental consistency. The models based on PERNN, PENN, PINN and FCNN all have a set of layers comprising a skip connections and batch normalization and a LeakyReLU activation function. Figure 17 illustrates the difference between conventional fully connected linear layers with activation function, and our skip-connection based layers. These layers involve the concatenation of earlier linear layers with the output of the activation layer (LeakyReLU) followed by a linear projection. We observed that these skip connection based layers allowed better convergence of the model, which can be seen by the results in Table 4 and explained in detail in section 5.2.3. To show the impact of these skip-connection based layers, we also include experiments on the versions of PERNN and PENN with conventional layers, labelled PERNN-NoSkip and PENN-NoSkip respectively.
Figure 17: A comparison of conventional and skip connection with batch normalization layers
In addition, Figure 18 shows the architectures of both PERNN and PINN models to illustrate the difference between the two approaches. In the PERNN model, the NEE ODE is a direct part of the computation graph inside the Physics Block, which outputs the final target variable ($\text{NEE}_{t+1}$). On the other hand, in the PINN model, the learnable layers directly predict the target NEE value ($\text{NEE}_{t+1}$) while the NEE ODE only offers external regularization to the loss function during training. In the PINN model, we present a branched architecture following similar structure to the learning blocks in PERNN. As seen in Figure 18, two of these branches outputs the parameters for the NEE ODE, while a dedicated branch outputs the target NEE value. During testing in inference mode, only this dedicated branch is active and predicts the target NEE value. This experiment allows us to present the impact of our knowledge blocks in the neural architecture with the Physics Block explicitly housing the NEE ODE as a direct part of the computational graph.
Figure 18: The architectures for the PERNN and PINN models. Top: PERNN architecture with two learning blocks, a residual block and physics block comprising NEE ODE as a direct part of computation graph. Bottom: PINN architecture with two branches outputting NEE ODE parameters which is used to regularize the loss function during training. A separate branch outputs the actual target NEE value.
5.2.3 Results
We evaluated the methods using three distribution-based metrics: Mean Maximum Discrepancy (MMD), Wasserstein Distance (Wsstn), and Kullback-Leibler Divergence (KL), to assess how well each technique captures the distribution of the target NEE. Additionally, we measured performance using mean absolute error (MAE) and R2 score to evaluate the fit to target variable. We also evaluate the accuracy of the intermediate variables from the learning blocks $E_{0}$, $r_{b,night}$ and $dT_{air}$, and the predicted time derivative of NEE, $d\text{NEE}$.
Table 4 presents the results on the test set over these metrics for the different methods compared in this study. It can be seen that PERNN surpasses the state-of-the-art Random Forest method by a notable 38.8% improvement in Wassertein Distance, 11% improvement in KL Divergence score while also achieving lower MAE score. PENN also shows comparable performance to the Random Forest albeit falling behind PERNN in all metrics, clearly indicating the advantage in convergence led by the incorporation of residual blocks in the physics encoded neural architecture. It is interesting to note that PINN performs considerably lower on all metrics compared to PERNN and PENN, which could potentially indicate the advantage of a direct integration of the NEE ODE in the computational graph of the neural network architecture on emprical grounds. Both PERNN and PENN also appear to outperform the FCNN and XGB methods on all metrics. It can also be seen that the PERNN-NoSkip and PENN-NoSkip perform considerably worse than their counterparts, which validates the benefit of the skip-connection based layers introduced in the learning and residual blocks of both PERNN and PENN.
Figure 19 shows the visualizations of ground-truth and predicted values for $\text{NEE}_{t+1}$ used for gap-filling across four time scales: daily, weekly, monthly and quarterly. Randomly sampled sequences, consistent across methods for fair validation, show that PERNN consistently captures NEE trends better than PENN, PINN, RF and XgBoost methods. It can also be seen that PENN, albeit falling short of PERNN, appears to capture these trends better than the rest of the methods in the experimentation.
Table 4: Results for NEE prediction on night time data and model experiments. The metrics MMD, Wasstn (Wasserstein Distance), KL (Kullback Leibler Divergence), and MAE (Mean Absolute Error) are expressed as the lower the better. R2 (score) is expressed as higher the better.
NEE $E_{0}$ $r_{b}$ $dT_{air}$ $d\text{NEE}$
Model MMD Wsstn KL MAE R2 MAE
PERNN 0.051 0.145 0.215 0.866 0.73 28.8 0.939 0.046 0.141
PERNN-NoSkip 0.138 0.396 0.260 1.087 0.67 29.9 1.29 0.036 0.11
PENN 0.086 0.190 0.212 0.986 0.69 30.7 1.66 0.08 0.063
PENN-NoSkip 0.122 0.276 1.149 1.17 0.48 31.06 1.51 0.014 0.053
PINN 0.247 1.043 0.540 1.41 0.55 27.33 1.826 0.029 0.031
FCNN 0.211 0.872 0.726 1.29 0.62 NA NA NA NA
RF 0.055 0.237 0.242 0.901 0.721 NA NA NA NA
XGB 0.052 0.202 0.214 0.988 0.658 NA NA NA NA
Figure 19: NEE predictions on test data across different time scales (daily, weekly, monthly, quarterly) for each approach in the experimentation : we create artificial gaps of the corresponding time scales that we fill with the corresponding model. The first column represents results for a single night in test set (2019-09-22), the second column represents results for a single week (2018-11-06 till 2018-11-11), the third columns represents results for a single month (2019-01-01 till 2019-01-29) and fourth column represents results for a single quarter (2019-01-01 till 2019-12-30). Row 1 shows results based on the PERNN model, Row 2 shows results based on the PENN model, Row 3 shows results from the PINN model, Row 4 shows results from the RF model, and Row 5 shows results from the XGB model respectively. The sequences illustrated in the graphs are randomly sampled from the test dataset and are kept consistent for each approach for fair validation.
@@ 6 Conclusion and Future Work
This paper presents a novel physics-encoded neural network architecture that integrates physics models directly into neural networks through residual knowledge blocks. This approach provides a general framework for incorporating differentiable physics models into data-driven learning algorithms, thereby enhancing reliability and interpretability by enabling the discovery of unobserved environmental variables. Furthermore, the proposed method offers a rapid prototyping capability for digital twin applications in scenarios where both observational data and detailed physical models are limited.
We evaluated our framework in two distinct application domains. The first application involved developing a steering model for autonomous vehicles in the TORCS simulation environment. In this context, our method outperformed pure data-driven fully-connected neural networks and traditional physics-informed neural networks by delivering improved generalizability on unseen road scenarios with significantly reduced model complexity and data requirements. The integration of differentiable geometrical and kinematic operators with learned intermediate variables—facilitated by residual and skip-connection based layers—proved critical for robust performance.
The second application demonstrated the versatility of our framework in a real-world digital twin setting for climate modeling. Specifically, we modeled Net Ecosystem Exchange (NEE) using an ordinary differential equation (ODE) derived from an Arrhenius-type formulation of ecosystem respiration. By embedding this ODE directly into the network via a dedicated physics block and training the model to predict key intermediate variables (e.g., temperature sensitivity $E_{0}$, base respiration $r_{b\_night}$, and the rate of change of ambient air temperature $\frac{dT}{dt}$), our method achieved superior gap-filling performance on flux tower data. Experimental results indicate that our approach outperforms state-of-the-art methods, such as Random Forest Robust and XGBoost, in both error metrics and the ability to capture the distributional characteristics of NEE.
In summary, our contributions demonstrate that directly incorporating physical models into neural network architectures not only enhances interpretability and prediction accuracy but also reduces reliance on extensive datasets. Future work will explore extending this framework to a broader array of digital twin applications where data scarcity and incomplete physics are prevalent. Additionally, further studies on gradient flow and the development of more robust residual blocks will be pursued to tackle increasingly complex physics-based models.
@@
  • Authors’ contributions: Muhammad Saad Zia is the primary author who contributed to the main ideation and experimentation in this manuscript. Corentin Houpert provided mathematical formulation for the Net Ecosystem Exchange ODE. Ashiq Anjum is the primary supervisor and reviewer on the work in this manuscript. Lu Liu is an academic reviewer and collaborator on the work in this manuscript. Anthony Conway and Anasol Peña-Rios are industrial collaborators on the work in this manuscript.